# Improving Intervention Efficacy via Concept Realignment in Concept Bottleneck Models

### v1

@@ Abstract
Concept Bottleneck Models (CBMs) ground image classification on human-understandable concepts to allow for interpretable model decisions. Crucially, the CBM design inherently allows for human interventions, in which expert users are given the ability to modify potentially misaligned concept choices to influence the decision behavior of the model in an interpretable fashion. However, existing approaches often require numerous human interventions per image to achieve strong performances, posing practical challenges in scenarios where obtaining human feedback is expensive. In this paper, we find that this is noticeably driven by an independent treatment of concepts during intervention, wherein a change of one concept does not influence the use of other ones in the model’s final decision. To address this issue, we introduce a trainable concept intervention realignment module, which leverages concept relations to realign concept assignments post-intervention. Across standard, real-world benchmarks, we find that concept realignment can significantly improve intervention efficacy; significantly reducing the number of interventions needed to reach a target classification performance or concept prediction accuracy. In addition, it easily integrates into existing concept-based architectures without requiring changes to the models themselves. This reduced cost of human-model collaboration is crucial to enhance the feasibility of CBMs in resource-constrained environments.
@@ 3.1 Background and Preliminaries
After each intervention $t$, we use a concept intervention policy $\pi(\tilde{c}_{t})$ to decide which concept to intervene on next. While $\pi$ can simply suggest random concepts for intervention, it is often much better to leverage heuristics that rank concepts in the order of importance (by some measure). A commonly deployed, effective intervention policy is UCP [18, 32], which utilizes the uncertainty of concepts. In particular, UCP selects concepts with the highest uncertainty, i.e. concept predictions closest to 0.5.
@@ 3.2 Concept Intervention Realignment
To address this, we propose a concept intervention realignment module (CIRM), which consists of two interdependent components: (a) a concept realignment model (CRM), $u:\mathcal{C}\rightarrow\mathcal{C}$. After a user intervenes on a subset of concepts $\mathcal{S}$, the remaining concepts ($\setminus\mathcal{S}$) are updated by a realigner network; and (b) an intervention policy $\pi$. The concepts predicted by the realignment model are fed to the policy to suggest which concept to intervene on next. Both components are interdependent, and together form the overall concept intervention realignment module, as also visualized in Figure 2. The training of the full CIRM comprising both selection policy and concept realignment model aims to simulate the complete intervention process. It thus starts from the concept predictions of the base model, $\hat{c}$, where we sequentially intervene on concepts for $T\leq k$ time steps by following a policy of choice, $\pi$ (in our case UCP by default, which we experimentally find to outperform random intervention significantly).
@@
Depending on the assumptions made on the realignment process, $v$ is either a simple MLP or a recurrent model (such as an LSTM [15]). The former parametrizes our default concept intervention realignment model, which only passes the set of intervened and un-intervened concepts at intervention step $t$ to the concept realignment model consisting of a simple MLP. The set of concepts fed into the MLP may either be the original concept embedding $\tilde{c}_{0}$, where all intervened concepts up to and including step $t$ have been replaced with ground-truth values, or the previously realigned $\kappa_{t-1}$ with similarly updated intervened concepts (c.f. Fig. 2, "GT"). Note that in either case, $\kappa_{t-1}$ informs the selection process of the subsequent concept to intervene on. After all interventions, the final concept embedding fed into the classifier is always $u(\tilde{c}_{T})$. Practically, we found using $\tilde{c}_{t}$ to work slightly better than $\kappa_{t-1}$. Both cases above however only pass the final set of concepts at time $t$ to the realignment model. Given the sequential nature of interventions, however, it may also be beneficial to account for the entire intervention history to inform future concept realignment. As a result, we also introduce a recurrent realignment variant, $u_{\text{rec}}$, which employs an LSTM model to retain the entire history of interventions until time $t$.
@@ 4.2 Concept Realignment Improves Intervention Efficacy
To probe the efficacy of our concept intervention realignment module, we evaluate both the change in concept prediction loss as well as overall classification accuracy as a function of intervened concept counts. These are visualized in Fig. 3 and Fig. 4 for sequentially trained CBMs (see §3.1), respectively, for all benchmark test sets - CUB, CelebA and AwA2. Note that for AwA2, we only show the first 50 interventions for visual clarity, as performance beyond that heavily plateaus since sufficient concepts have been intervened on to perfectly solve the test data. Table 1 numerically summarizes these results via AUC scores and provides additional scores for independently trained CBMs, jointly trained CBMs as well as Concept Embedding Models (CEMs). Runs in Tab. 1 and Figs. 34 all utilize the stronger UCP concept selection policy as opposed to the weaker random selection policy to measure intervention efficacy at the highest level, and train the concept realignment module on top of already trained concept models.
@@
Across all datasets, we can observe a consistent, in parts vast reduction in concept prediction loss, which measures the correct assignment of concepts for each input (using the concept loss described in §3.1). For example on CUB, a tenfold reduction of the original unintervened concept loss ($\sim{}0.6$ to $\sim{}0.06$) can be achieved with half the number of interventions ($11$ with concept realignment, $23$ without). This effect becomes even more prevalent on AwA2, where a tenfold reduction ($017$ 0.017) is achieved after around $16$ interventions with realignment versus more than $60$ without; marking a more than $70\%$ reduction in intervention efforts. This is also reflected in Tab. 1, where concept loss AUC drops by in parts more than half for CUB and from $4.26$ to $1.13$ on AwA2. We find this significant improvement in concept attribution persists across all CBMs and CEMs.
@@ 4.3 Intervention Realignment for Intervention-aware CEMs
Our results are shown in Fig. 5. First, we find that explicit concept intervention realignment can significantly improve correct concept attribution, even in intervention-aware training setups (c.f. Fig. 5a). While not as significant as improvements over standard CBM models, for specific target concept prediction losses (such as a fivefold reduction from $0.5$ to $0.1$), half the number of intervention steps are needed ($11$ versus $20$). The improved concept attribution is also reflected in higher intervention accuracies as seen in Fig. 5b, albeit the overall (still notable!) improvement is less reflective of the significant gains on a concept level. Overall, however, our experiments highlight that even when applied to state-of-the-art approaches that specifically simulate the intervention process during training, improved intervention efficacy can be found. Importantly, the consistently significant improvements on a concept attribution level mean that classification decisions are much better grounded on correct concept attributions, which is crucial for interpretability [17, 40] of classification results. Finally, we find that concept intervention realignment can be applied both as a regularization mechanism during training, as well as adapted entirely posthoc, while still offering consistent benefits. This supports the high versatility of CIRMs as a general-purpose tool to increase intervention efficacy.

### v2

@@ Abstract
Concept Bottleneck Models (CBMs) ground image classification on human-understandable concepts to allow for interpretable model decisions. Crucially, the CBM design inherently allows for human interventions, in which expert users are given the ability to modify potentially misaligned concept choices to influence the decision behavior of the model in an interpretable fashion. However, existing approaches often require numerous human interventions per image to achieve strong performances, posing practical challenges in scenarios where obtaining human feedback is expensive. In this paper, we find that this is noticeably driven by an independent treatment of concepts during intervention, wherein a change of one concept does not influence the use of other ones in the model’s final decision. To address this issue, we introduce a trainable concept intervention realignment module, which leverages concept relations to realign concept assignments post-intervention. Across standard, real-world benchmarks, we find that concept realignment can significantly improve intervention efficacy; significantly reducing the number of interventions needed to reach a target classification performance or concept prediction accuracy. In addition, it easily integrates into existing concept-based architectures without requiring changes to the models themselves. This reduced cost of human-model collaboration is crucial to enhance the feasibility of CBMs in resource-constrained environments. Our code is available at https://github.com/ExplainableML/concept_realignment.
@@ 3.1 Background and Preliminaries
After each intervention $t$, we use a concept intervention policy $\pi(\tilde{c}_{t})$ to decide which concept to intervene on next. While $\pi$ can simply suggest random concepts for intervention, it is often much better to leverage heuristics that rank concepts in the order of importance (by some measure). A commonly deployed, effective intervention policy is UCP [18, 32], which utilizes the uncertainty of concepts. In particular, UCP selects concepts with the highest uncertainty, i.e. concept predictions closest to 0.5. More details about the intervention process can be found in Algorithm 1.
@@ 3.2 Concept Intervention Realignment
To address this, we propose a concept intervention realignment module (CIRM), which consists of two interdependent components: (a) a concept realignment model (CRM), $u:\mathcal{C}\rightarrow\mathcal{C}$. After a user intervenes on a subset of concepts $\mathcal{S}$, the remaining concepts ($\setminus\mathcal{S}$) are updated by a realigner network; and (b) an intervention policy $\pi$. The concepts predicted by the realignment model are fed to the policy to suggest which concept to intervene on next. Both components are interdependent, and together form the overall concept intervention realignment module, as also visualized in Figure 2. The training of the full CIRM comprising both selection policy and concept realignment model aims to simulate the complete intervention process. It thus starts from the concept predictions of the base model, $\hat{c}$, where we sequentially intervene on concepts for $T\leq k$ time steps by following a policy of choice, $\pi$ (in our case UCP by default, which we experimentally find to outperform random intervention significantly; See Supp. §B).
@@
Depending on the assumptions made on the realignment process, $v$ is either a simple MLP or a recurrent model (such as an LSTM [15]). The former parametrizes our default concept intervention realignment model, which only passes the set of intervened and un-intervened concepts at intervention step $t$ to the concept realignment model consisting of a simple MLP. The set of concepts fed into the MLP may either be the original concept embedding $\tilde{c}_{0}$, where all intervened concepts up to and including step $t$ have been replaced with ground-truth values, or the previously realigned $\kappa_{t-1}$ with similarly updated intervened concepts (c.f. Fig. 2, "GT"). Note that in either case, $\kappa_{t-1}$ informs the selection process of the subsequent concept to intervene on. After all interventions, the final concept embedding fed into the classifier is always $u(\tilde{c}_{T})$. Practically, we found using $\tilde{c}_{t}$ to work slightly better than $\kappa_{t-1}$. Both cases above however only pass the final set of concepts at time $t$ to the realignment model. Given the sequential nature of interventions, however, it may also be beneficial to account for the entire intervention history to inform future concept realignment. As a result, we also introduce a recurrent realignment variant, $u_{\text{rec}}$, which employs an LSTM model to retain the entire history of interventions until time $t$. An algorithmic summary is provided in supplementary §A.
@@ 4.2 Concept Realignment Improves Intervention Efficacy
To probe the efficacy of our concept intervention realignment module, we evaluate both the change in concept prediction loss as well as overall classification accuracy as a function of intervened concept counts. These are visualized in Fig. 3 and Fig. 4 for sequentially trained CBMs (see §3.1), respectively, for all benchmark test sets - CUB, CelebA and AwA2. Note that for AwA2, we only show the first 50 interventions for visual clarity, as performance beyond that heavily plateaus since sufficient concepts have been intervened on to perfectly solve the test data. Table 1 numerically summarizes these results via AUC scores and provides additional scores for independently trained CBMs, jointly trained CBMs as well as Concept Embedding Models (CEMs). Runs in Tab. 1 and Figs. 34 all utilize the stronger UCP concept selection policy as opposed to the weaker random selection policy (§B) to measure intervention efficacy at the highest level, and train the concept realignment module on top of already trained concept models.
@@
Across all datasets, we can observe a consistent, in parts vast reduction in concept prediction loss, which measures the correct assignment of concepts for each input (using the concept loss described in §3.1). For example on CUB, a tenfold reduction of the original unintervened concept loss ($\sim{}0.6$ to $\sim{}0.06$) can be achieved with half the number of interventions ($11$ with concept realignment, $23$ without). This effect becomes even more prevalent on AwA2, where a tenfold reduction ($017$ 0.017) is achieved after around $16$ interventions with realignment versus more than $60$ without; marking a more than $70\%$ reduction in intervention efforts. This is also reflected in Tab. 1, where concept loss AUC drops by in parts more than half for CUB and from $4.26$ to $1.13$ on AwA2. We find this significant improvement in concept attribution persists across all CBMs and CEMs, as well as random seed initializations (see Supp. Tab. 2,  3 and  4)
@@ 4.3 Intervention Realignment for Intervention-aware CEMs
Our results are shown in Fig. 5. First, we find that explicit concept intervention realignment can significantly improve correct concept attribution, even in intervention-aware training setups (c.f. Fig. 5a). While not as significant as improvements over standard CBM models, for specific target concept prediction losses (such as a fivefold reduction from $0.5$ to $0.1$), half the number of intervention steps are needed ($11$ versus $20$). The improved concept attribution is also reflected in higher intervention accuracies as seen in Fig. 5b, albeit the overall (still notable!) improvement is less reflective of the significant gains on a concept level (additional results can be found in Supp. §C). Overall, however, our experiments highlight that even when applied to state-of-the-art approaches that specifically simulate the intervention process during training, improved intervention efficacy can be found. Importantly, the consistently significant improvements on a concept attribution level mean that classification decisions are much better grounded on correct concept attributions, which is crucial for interpretability [17, 40] of classification results. Finally, we find that concept intervention realignment can be applied both as a regularization mechanism during training, as well as adapted entirely posthoc, while still offering consistent benefits. This supports the high versatility of CIRMs as a general-purpose tool to increase intervention efficacy.
@@ Acknowledgements
Appendix A Details of the Intervention Procedure
In Algorithm 1 we describe the standard process of performing interventions in concept-based models using an intervention policy.
Algorithm 1 Intervention Algorithm
1:Inputs:
2:      $T$ (total number of interventions)
3:      $\pi$ (intervention policy, which takes the concepts as input)
4:      $\hat{c}$ (concepts predicted by the concept encoder)
5:$\tilde{c}\leftarrow\hat{c}$triangleright output of the concept encoder, g
6:for $t\in\{0,\ldots,T-1\}$ do
7:     $i\leftarrow\pi(\tilde{c})$triangleright $i$ is the concept that we want the user to intervene on
8:     $\tilde{c}_{i}\leftarrow c_{i}$triangleright replace the $i$th concept in $\tilde{c}$ with its ground truth value $c_{i}$
9:end for
10:return $\tilde{y}=f(\tilde{c})$triangleright updated class prediction after all interventions have been performed
In Algorithm 2 we describe the procedure used in our setup, which realigns unintervened concepts following an intervention step. We use this algorithm to compute the loss for training the realignment model.
Algorithm 2 Realignment Model Training Loss
1:Inputs:
2:      $T$ (total number of interventions)
3:      $\pi$ (intervention policy, which takes the concepts as input)
4:      $\hat{c}$ (concepts predicted by the concept encoder)
5:$\tilde{c}\leftarrow\hat{c}$triangleright output of the concept encoder, g
6:$\kappa_{-1}\leftarrow\hat{c}$triangleright initialize realigned concepts
7:$\mathcal{L}\leftarrow 0$triangleright initialize loss
8:for $t\in\{0,\ldots,T-1\}$ do
9:     $i\leftarrow\pi(\kappa_{t-1})$triangleright $i$ is the concept that we want the user to intervene on
10:     $\tilde{c}_{i}\leftarrow c_{i}$triangleright replace the $i$th concept in $\tilde{c}$ with its ground truth value $c_{i}$
11:     $\kappa_{t}\leftarrow u(\tilde{c})$triangleright output of realignment model
12:     $\mathcal{L}\leftarrow\mathcal{L}+\textrm{CE}(\kappa_{t},c)$triangleright aggregate loss
13:end for
14:return $\mathcal{L}/T$triangleright average loss across all intervention steps
Appendix B Comparison Between Random and UCP Policies
In this section, we compare the classification accuracies achieved by following the random and UCP intervention policies on the three datasets, respectively. In Fig. 10 we show that the UCP policy is superior across all datasets, and is therefore our default policy across all experiments in this study.
Figure 10: Comparison between accuracy under UCP and Random intervention policies. UCP is superior in all three datasets.
Appendix C Additional Results on IntCEMs
In this section, we report the performance of posthoc concept realignment on the intervention-aware CEMs (IntCEMs) on the CelebA and AwA2 datasets to supplement the results in Section 4.3. In Fig. 11 we show that concept realignment improves the performance of the SoTA approach in both datasets.
Figure 11: Classification accuracy with and without posthoc concept realignment in intervention-aware CEMs. In both cases, concept realignment improves performance of the base IntCEM model.
Appendix D Additional Results
Here, we report the area under the curve of concept prediction loss and classification accuracy using three different random seeds. It can be seen in Tables 2 and 3 that concept realignment consistently improves performance on both metrics.
Table 2: Area Under Curve (AUC) of Concept Prediction Loss and Classification Accuracy with/without CIRM for three random seeds on the CUB dataset. We use the same backbone for sequential and independent CBMs. CIRM improves performance across all models and runs.
Run 1 Run 2 Run 3 Run 1 Run 2 Run 3
[…]
Table 3: Area Under Curve (AUC) of Concept Prediction Loss and Classification Accuracy with/without CIRM for three random seeds on the CelebA dataset. We use the same backbone for sequential and independent CBMs. CIRM improves performance across all models and runs.
Run 1 Run 2 Run 3 Run 1 Run 2 Run 3
[…]
Table 4: Area Under Curve (AUC) of Concept Prediction Loss and Classification Accuracy with/without CIRM for three random seeds on the AwA2 dataset. We use the same backbone for sequential and independent CBMs. CIRM improves performance across all models and runs.
Run 1 Run 2 Run 3 Run 1 Run 2 Run 3
[…]